# Unveiling Molecular Moieties through Hierarchical Graph Explainability

### 29 Jan 2024 v1

@@ Abstract
Background: Graph Neural Networks (GNN) have emerged in very recent years as a powerful tool for supporting in silico Virtual Screening. In this work we present a GNN which uses Graph Convolutional architectures to achieve very accurate multi-target screening. We also devised a hierarchical Explainable Artificial Intelligence (XAI) technique to catch information directly at atom, ring, and whole molecule level by leveraging the message passing mechanism. In this way, we find the most relevant moieties involved in bioactivity prediction.
@@
Conclusion: Our approach is a valid support for shortening both the screening and the hit-to-lead phase. Detailed knowledge about the molecular substructures that play a role in the inhibitory action, can help the computational chemist to gain insights into the pharmacophoric function of the molecule also for repurposing purposes.
@@ 1 Background
The development of a new drug is a long and time-consuming process that, despite advances in technology and computer applications, is fragmented into several steps before being able to identify a lead compound that can be tested on in vivo models. Over the last decade, the use of deep learning has provided support to the Virtual Screening (VS) campaigns for the efficient selection of hit compounds by strengthening predictions through the use of massive data in the training model phase [11, 4, 22, 3, 29]. Indeed, being able to correctly prioritise active molecules during the screening steps speeds up the discovery process. Classical approaches used in CADD are not always efficient, being affected by the use of specific force field [21, 13] and the quality of structural data available, thus affecting the hit optimisation steps towards the lead compound. Hit discovery is the very first step after screening, and chemists use it as the starting point to optimise hits rationally until designing a lead compound. This step uses different techniques such as hit expansion, (bio)isosteric replacement or a combination of them [10]. Hit expansion, for instance, is one of the most widely used techniques for optimisation and it is based on the synthesis of analogues to the selected small molecule by replacing chemical patterns, guided by data obtained from Structure Activity Relationship (SAR) analysis [2]. (Bio)isosteric replacement is a technique based on the definition of isoterism introduced by Langmuir in 1919, during his studies of the physicochemical properties of atoms and chemical groups. He noted that atoms with similar properties trigger the same biological effect; for example, compounds with the same number of atoms and electrons (e.g. N, CO, NO, etc.) are isosteric and trigger the same biological behaviour. On the basis of the same electron arrays, the chemical groups have been divided into 21 groups that are used to optimise bioactive molecules by increasing their efficiency as [7] drugs. Clearly, optimisation requires an understanding of the relevant atoms or chemical groups within the small molecule. This step is usually handled by the computational chemist, who recognizes and consciously acts to improve the molecule’s activity in accordance with their knowledge and expertise in the field.
@@
In the present work, we are presenting a GNN based on the so called Graph Convolutional layers, which firstly prioritise molecules towards a set of 20 kinases with a very high rate of precision, and also supports the computational chemist in the selection of the most important chemical features in the host-guest recognition. The proposed neural architecture provides a very good bioactivity classification (F score 98% on average). Moreover, we devised a suitable Explainable AI (XAI) procedure, which is inspired to the well known Gradient-weighted Class Activation Mapping (Grad-CAM) [19]. Grad-CAM is an explainer used in image classification through Convolutional Neural Networks (CNN) and it generates a heat-map for each class to highlight the parts of an image that contribute to the recognition of that class. At each layer where Grad-CAM is applied, the role of each feature map in explaining the class is weighted using the gradients back propagated to the layer itself. In [16] a graph based version of Grad-CAM is proposed, and we used this implementation for our molecular explainer that provides a graphical representation highlighing the atoms and/or chemical groups with a key role in the discriminative event. We used the four macro-groups of pharmacophoric features (H-bond donor or acceptor, hydrophobic, aromatic) [6] in this respect.
@@ 2.1 Data preparation
Canonical SMILES were recovered from ChEMBL [1] for all the small molecules whose activity on the target protein is referenced by IC, K, and K values. We used a fixed threshold activity value on each target, in order to generate the molecular graphs through the use of the RDKit library, which allows obtaining all the features previously described. The final dataset consists of 89373 small molecules and has a 1:100 active/inactive ratio for the smallest class that is CLK2. This ratio allows profound assessing of the model generalization abilities as it has more or less the same size of a real world screening task. Training, validation, and test sets were created using a 80%:10%:10% ratio for each target, in order to minimize the comparison bias.
@@ 2.2 Graph Convolutional Neural Network
The message-passing algorithm stacks data coming from neighbours, for each node in the graph. This aggregation process results in an updated set of node features placed along the channel dimension of the data sample. More formally, the algorithm uses a message function and a vertex update function , which is usually implemented as a Multi Layer Perceptron (MLP). Both those functions are learnable and differentiable in order to be trained in a neural network. Given the time-step , the node collecting features, the edge connecting node and any node in the set of the neighbors of in the graph , and the hidden states and , the new message at time-step is calculated as shown in eq. 1.
The new hidden state is then computed using the previous hidden state and the new message , shown in eq. 2.
@@ 2.3 Explainability analysis
Explainable AI (XAI) techniques were chosen to explore the black box of operations underlying neural network prediction. In the proposed work, XAI algorithms were used to be able to identify the parts of the small molecule directly involved in bioactivity. In fact, each hit molecule consists of groups with relevant activity in the interaction with the target that can be labelled as pharmacophoric.
Two different explainability tools were considered for the evaluation of the network results. Both are part of the so-called ”Instance-level methods”, i.e. those that provide a solution by identifying the most important input features that contribute most to the prediction. One of them, GNNExplainer [27], was picked because it is a ”perturbation-based method” that tracks the change in output for minor variations of the same input data to determine which features are most important for prediction. However, this approach simply offers the graph arcs’ contributions to the label prediction. Due to this, a second technique from the class of ”Gradient and Feature-based techniques” was selected. With this method, a high gradient or higher values in the feature maps suggest that such features have a larger importance for the predictions. The goal is to use the gradient or hidden feature maps of the model to show the importance of input features.
An implementation from the work of Pope et Al. [16] was utilized for this purpose. In this work, the authors implemented some graph-based counterparts of three well-established explainability methods originally designed for convolutional neural networks. Among those, Gradient-weighted Class Activaction Mapping (GradCAM) [19] has been chosen as the most representative of relevant input features for prediction. This method aims to visualize the region that activate most strongly when the network makes prediction for a particular class. It is done by examining the gradients of the target class score with respect to the feature map in the network.
The formula for calculating the weights , related to layer , node and number of channels of the feature maps , is shown in eq. 3.
Grad-CAM values for each node is then calculated as described here in eq. 4.
The original Grad-CAM implementation has been modified to be optimized and suited for this task. First of all, unlike from the original implementation, the ReLU activation function has been removed. The decision was made to capture not only positive contributions to the prediction using Grad-CAM values but to include contributions from both negative and positive influences. Far from the image-related task of the original implementation, our explainability goal is to get information about pharmacophoric features of the input graph and not just insights about biological activity of the molecule.
The analysis of the features importance has been conducted performing the GradCAM evaluation multiple times, each time referring to a different ”level” of depth of the convolutional segment of the network. In a graph convolutional neural network, the message-passing algorithm is crucial for the information retrieval. For each step of the graph convolution layer, a message representing knowledge about neighbours is aggregated into the channel dimension of each node. In this scenario of molecule activity prediction we first chose to use GradCAM to look at those contribution after the first round of graph convolutions, for considering the atoms’ impact for a 1-step neighbourhood. In the same way, we chose to look after the third and the seventh (last) convolutional operation. In this second case, we want to highlight contribution for a 3-step neighbourhood, focusing on chemical structures that could be identified with a 3-step path for each of the atoms, such as rings. Focusing on the output of the last convolution, we make sure that we can have contribution for a 7-step path starting from each of the atoms in the molecule.
@@ 3.1 Multi-class multi-target classification
In this section results from the 20-target classifier are discussed. Two accuracy metrics are used to describe the network’s performance: first, Global Balanced Accuracy assesses the network’s ability to predict both active and inactive molecules, while Sensitivity (also named Recall) focuses solely on the network’s performance in correctly classifying active molecules. The F1-score quantifies how well the model can correctly identify positive instances (Recall) while avoiding false positives (Precision). On the other hand, the ROC-AUC curve measures the model’s ability to discriminate between positive and negative instances. The TP/P ratio is employed to assess the classifier’s ability to prioritize ligands, counting the correct predictions relative to the total positives in the top x% of the test set. Finally, the Enrichment Factor metric measures how many times the tested model outperforms a purely random process. Both TP/P and EF were calculated for 1, 2, 5 and 10% of the whole test set.
@@
The findings show that the network correctly identifies positive and negative samples in terms of balanced accuracy and sensitivity. The range of results for this last metric increased significantly from the earlier work [14], where it ranged between 0.22 and 0.84. Now sensitivity has an average value of 0.904, showing that it is much easier for the network to classify positive samples. F1-Score results show that the network is able to overcome the dataset imbalance while simultaneously reducing the number of false negatives and false positives. The use of a weighted cross entropy loss function and a soft F-score loss term each properly balanced worked fine for the classification task.
Results for metrics TP/P and EF are shown in Table 2. The percentage of test set considered for 1%, 2%, 5% and 10% are respectively 134, 268, 670 and 1340. Both TP/P and EF are close to the ones of the previous work [14].
Results in table 3 refer to the single target classifier performances on the CDK1 protein.
@@
The results obtained with the multiclassifier are excellent in terms of sensitivity, outperforming the multiclassifier based on the embedding of molecular fingerprints proposed by the same authors in the paper Mendolia et al. [14] The comparison between the two results is summarised in Figure 3 showing the sensitivity for all targets. It is notable that especially for the case of CLK2 there was a consistent improvement in performance related to sensitivity metric, going from a value of (—) to a value of 0.938, as it is shown also in Table 1.
@@ 3.2 Single-target classification
The multi-target classification architecture was not suited to the explainability applications proposed by the authors. In fact, in order to be able to accurately identify the pharmacophoric groups that characterise high-priority molecules, a model trained on CDK1 has been created to reduce the bias provided by the different targets. Classification performance was assessed using the same metrics as described in the 3.1 section and is shown in table 3 and 4.
@@
| CDK1 | 34/109 | 58/109 | 109/109 | 100/109 | 31.19 | 26.60 | 2 | 9.11 |
As can be observed, both Global Balanced accuracy and Sensitivity are higher than those obtained on the same target in the multiclassifier, while performance in terms of TP/P and EF remained consistent with that observed previously.
@@ 3.3 Explainability results
The XAI GRAD-CAM-like algorithm has been used to identify the molecular regions with the higher impact on the target assignment. As described in the section 2.3 the evaluation of gradients at three different depths was the main insight of the work presented. In fact, the message-passing mechanism aggregates node information at each GraphConvolution. In this way, the first level gathers information on individual atoms, the second level integrates the information with that of neighbouring nodes up to a maximum of 3 nodes, in order to identify functional groups, while with the last level the entire structure is taken into account, bringing out entire regions, commonly called molecular substructures, that actively contribute to the classification. An example of what is observed with the gradient analysis for the different depths is shown in figure 4, where the parts of the molecules with higher contribute are highlighted.
The figure 4 shows the atoms or their atomic groups which are identified as relevant to the 3 depths of explanation. Of particular interest is also the chemical interpretation attributed to each group, in fact the 3 different pharmacophoric categories and a fourth category within which all the relevant atoms will be placed. In fact, the H-bond acceptor or donor groups are highlighted in yellow, the hydrophobic groups in blue, and the aromatic groups in green. Lastly, the relevant atoms, which do not have explicit pharmacophoric functions but whose contribution in terms of bioactivity and/or stabilisation of the interaction with the binding site is highlighted, are highlighted in red.
The explanation was then tried on 19 CDK1 inhibitors, and the results has been compared to the docking results or experimental crystal structures of co-crystalised ligands (when available) for the same inhibitors. For 17 of 19 known drugs, our methodology is consistent with what was observed using experimental structures. and Table 1 shows which drug interactions were predicted by the proposed explainer. Figures 5 and 6 show an example of the information gathered during this phase of testing, which is shown in the supplementary material for all tested drugs. Interestngly, every ”level” of depth of the convolution segment gives an important contribution to the final molecular interpretation. Indeed, as observed for Dinacilcib (Fig.1) the second step of depth (the one considering three atoms connected) is capable to catch important features as the H-bond donor N6, useful for the ligand protein complex stabilisation. Such a crucial information is then lost in the next step of the ligand interpretation (level7) where molecular substructures are considered. On the other hand, the higher level of molecular interpretation, is crucial to catch all the aromatic and lipophilic moieties that could be lost in the previous stages of graph representation.
| Drug | Pubchem ID | GFP* | Accordance to docking | ||||
@@
| C26H31N7O** | 46916588 | yes |
*Guest Feature Prediction
**Molecular formula of 2-[(2S)-1-[8-ethyl-4-[(4-pyridin-2-ylphenyl)methylamino]pyrazolo[1,5-a][1,3,5]triazin-2-yl]piperidin-2-yl]ethanol
| Drug | 3 level XAI | 7 level XAI | Accordance to docking | ||||||||||||
@@
*Guest Feature Prediction
**Molecular formula of 2-[(2S)-1-[8-ethyl-4-[(4-pyridin-2-ylphenyl)methylamino]pyrazolo[1,5-a][1,3,5]triazin-2-yl]piperidin-2-yl]ethanol

### 29 Feb 2024 v2

@@ Abstract
Background: Graph Neural Networks (GNN) have emerged in very recent years as a powerful tool for supporting in silico Virtual Screening. In this work we present a GNN which uses Graph Convolutional (GC) architectures to achieve very accurate multi-target screening. We also devised a hierarchical Explainable Artificial Intelligence (XAI) technique to catch information directly at atom, ring, and whole molecule level by leveraging the message passing mechanism. In this way, we find the most relevant moieties involved in bioactivity prediction.
@@
Conclusion: Our approach is a valid support for shortening both the screening and the hit-to-lead phase. Detailed knowledge about the molecular substructures that play a role in the inhibitory action, can help the computational chemist to gain insights into the pharmacophoric function of the molecule also for repurposing tasks.
@@ 1 Background
The development of a new drug is a long and time-consuming process that, despite advances in technology and computer applications, is fragmented into several steps before being able to identify a lead compound that can be tested on in vivo models. Over the last decade, the use of deep learning has provided support to the Virtual Screening (VS) campaigns for the efficient selection of hit compounds by strengthening predictions through the use of massive data in the training model phase [11, 4, 22, 3, 29]. Indeed, being able to correctly prioritise active molecules during the screening steps speeds up the discovery process. Classical approaches used in Computer Aided Drug Design (CADD) are not always efficient, being affected by the use of specific force field [21, 13] and the quality of structural data available, thus affecting the hit optimisation steps towards the lead compound. Hit discovery is the very first step after screening, and chemists use it as the starting point to optimise hits rationally until designing a lead compound. This step uses different techniques such as hit expansion, (bio)isosteric replacement or a combination of them [10]. Hit expansion, for instance, is one of the most widely used techniques for optimisation and it is based on the synthesis of analogues to the selected small molecule by replacing chemical patterns, guided by data obtained from Structure Activity Relationship (SAR) analysis [2]. (Bio)isosteric replacement is a technique based on the definition of isosterism introduced by Langmuir in 1919, during his studies of the physicochemical properties of atoms and chemical groups. He noted that atoms with similar properties trigger the same biological effect; for example, compounds with the same number of atoms and electrons (e.g. N, CO, NO, etc.) are isosteric and trigger the same biological behaviour. On the basis of the same electron arrays, the chemical groups have been divided into 21 groups that are used to optimise bioactive molecules by increasing their efficiency as [7] drugs. Clearly, optimisation requires an understanding of the relevant atoms or chemical groups within the small molecule. This step is usually handled by the computational chemist, who recognizes and consciously acts to improve the molecule’s activity in accordance with their knowledge and expertise in the field.
@@
In the present work, we are presenting a GNN based on the so called Graph Convolutional layers, which firstly prioritise molecules towards a set of 20 kinases with a very high rate of precision, and also supports the computational chemist in the selection of the most important chemical features in the host-guest recognition. The proposed neural architecture provides a very good bioactivity classification (F score 98% on average). Moreover, we devised a suitable Explainable AI (XAI) procedure, which is inspired to the well known Gradient-weighted Class Activation Mapping (Grad-CAM) [19]. Grad-CAM is an explainer used in image classification through Convolutional Neural Networks (CNN) and it generates a heat-map for each class to highlight the parts of an image that contribute to the recognition of that class. At each layer where Grad-CAM is applied, the role of each feature map in explaining the class is weighted using the gradients backpropagated to the layer itself. In [16] a graph based version of Grad-CAM is proposed, and we used this implementation for our molecular explainer that provides a graphical representation highlighting the atoms and/or chemical groups with a key role in the discriminative event. We used the four macro-groups of pharmacophoric features (H-bond donor or acceptor, hydrophobic, aromatic) [6] in this respect.
@@ 2.1 Data preparation
Canonical SMILES were recovered from ChEMBL [1] for all the small molecules whose activity on the target protein is referenced by IC, K, and K values. We used a fixed threshold activity value on each target, in order to generate the molecular graphs through the use of the RDKit library, which allows obtaining all the features previously described.
The final dataset consists of 89373 small molecules and has a 1:100 active/inactive ratio for the smallest class that is CLK2. This ratio allows profound assessing of the model generalization abilities as it has more or less the same size of a real world screening task. Training, validation, and test sets were created using a 80%:10%:10% ratio for each target, in order to minimize the comparison bias.
@@ 2.2 Graph Convolutional Neural Network
The message-passing algorithm stacks data coming from neighbours, for each node in the graph. This aggregation process results in an updated set of node features placed along the channel dimension of the data sample. More formally, the algorithm uses a message function and a vertex update function , which is usually implemented as a Multi Layer Perceptron (MLP). Both those functions are learnable and differentiable in order to be trained in a neural network. Given the time-step , the node collecting features, the edge connecting node and any node in the set of the neighbors of in the graph , and the hidden states and , the new message at time-step is calculated as:
The new hidden state is then computed using the previous hidden state and the new message using the vertex update function:
@@ 2.3 Explainability analysis
As already mentioned, we used XAI techniques to gain information about the most relevant molecular moieties involved in bioactivity prediction; different algorithms were tried in this respect. All the selected XAI approaches belong to the so-called “instance-level methods” i.e. the ones aimed at identifying the input features that contribute mostly to the prediction. At first, we tried GNNExplainer [27], which is a “perturbation-based method” that tracks the change in the output for minor variations of the same input data to determine which features are most important for prediction. GNNExplainer provides a too fine-grained information because it works at feature level, while we were interested to obtain explanations at least at atom (i.e. node) level. As a consequence, we moved towards the class of “gradient and feature-based techniques”. An implementation from the work by Pope et al. [16] was utilized for this purpose. In this work, the authors implemented some graph-based counterparts of three well-established explainability methods, which had been designed originally for convolutional neural networks.
We choose Grad-CAM as the best performing technique in labeling point-wise contributions to positive prediction, at least in the original image domain where pixels are the data units, and the information flows between layers through the convolution operator. In a GC layer, feature maps are built for each node of the graph by message passing (eq. 1). In this way, features at each node encode the information flow from the node neighborhood, and the more deep is the layer the more its features collect information from many message passing steps computed in the previous layers (eq. 2). As a consequence, a feature map at a certain depth collects information from the neighbors of a node.
Grad-CAM aims at visualizing the most active region of the input, when a particular class is predicted. This is done by examining the gradients of the target class score with respect to each feature map in the network layer under investigation. The weight , for class related to the -th node feature at layer is computed as:
is the -th feature value for the single node in layer , while is the score for class obtained from the last network layer, before the output unit. Obviously, the partial derivative term is computed using the chain rule when is not the last layer. Grad-CAM values for the whole node are summed over its features, and passed to the ReLU activation function to filter out negative contributions to the prediction of class .
In our implementation that is we have just a single class score because we implemented our explainer on a binary classifier. Moreover, the ReLU activation function has been removed to capture not only positive contributions to the prediction, but also negative ones. Our choice is motivated by the intrinsic nature of the classification task to be explained. In the image domain, either single pixels are classified as belonging to a particular class, in the case of semantic segmentation, or the whole image is labelled according to the presence or not of some class discriminative features that are learned by the network. In any case, it is possible to think about classification as a pattern recognition task. As a consequence, the explainer tends to isolate the regions that contributed to the prediction, while discarding the others, and the ReLU activation function is intended for this purpose.
Bioactivity prediction is a classification task where molecules are labelled as positive or negative if they trigger “globally” a particular chemical event or not. Many factors are involved in bioactivity apart from the presence of the classical pharmacophoric structures: molecular chirality, hydrosolubility, use of conformational restriction in the drug design, and so on. All of them can be captured by training the GCNN on a three-dimensional molecular representation with rich information at the nodes, but can not be directly related to just “positive” or “negative” Grad-CAM values at each node. As a consequence we removed ReLU for our explainability task. Moreover, pharmacophoric information is strictly related to the presence in the molecule of particular structures that span atoms’ neighborhoods with different size. In view of this consideration we evaluated Grad-CAM multiple times at different depths in the convolutional segment of the network. We used the classical Grad-CAM arrangement after the last layer of the network that is the 7-th one. This explainer is aimed at highlighting relevant features globally, and it is not able to provide information about subtle moieties like acceptors or donors. To reach this aim, we placed also a Grad-CAM explainer at the very first GC layer to consider the atoms’ impact in a 1-neighbourhood that is at a bond level. Another explainer was placed after the third layer to catch information related to 3-neighborhoods. At this level, as we analyze the 3-neighborhood of each atom, the action of an entire substructure like aromatic ring or carbonyl ring can be detected.
The whole explanation map has been built as follows. At first, the three Grad-CAM values distribution have been summed for each node:
The resulting distribution has been min-max scaled, and only contributions higher than have been regarded as relevant. The RDKit††https://www.rdkit.org/ software package has been used to devise pharmacophoric labels for each molecular structure, and the final map reports only relevant contributions colored accordingly to the labels. In particular, H-bond acceptor or donor groups are highlighted in yellow, hydrophobic groups are colored in blue, and the aromatic groups are green. Lastly, atoms without an explicit pharmacophoric label where are labelled as “relevant”, and are colored in red. In Section LABEL:exp-results a detailed analysis is reported about the structures highlighted by our hierarchical Grad-CAM explainer.
@@ 3 Results
In this section we present the results for both the multi-target and the CDK1-only classifier, and the explainability analysis is reported for nineteen anti-cancer drugs that are CDK1 inhibitors.
@@ 3.1 Multi-class multi-target classification
Results for the 20-target classifier are shown in Table 1 and Table 2. The first table provides the performance metrics of our network as a classifier. The Balanced Accuracy assesses the network’s ability yo make correct predictions despite the class imbalance, while the Sensitivity (also named Recall) focuses solely on the network’s performance in correctly classifying active molecules. The F1-score quantifies how well the model can correctly identify positive instances (Recall) while avoiding false positives (Precision). On the other hand, the Area Under the Curve (AUC) provides a global measure of the model’s ability to discriminate between positive and negative instances, and it is computed from the curve that plots the Sensitivity vs the False Positive Rate varying the threshold to discriminate positives and negatives.
The second Table reports the measure used to assess the classifier’s ability to prioritize ligands, that is the true outcome of the VS task. The True Positives vs Positives ratio (TP/P) is employed to assess correct predictions count relative to the total positives in a given class, the Enrichment Factor (EF) metric measures how many times the tested model outperforms a purely random process in predicting a given class. Both these measures are computed for a given topmost percentage of the test set, and we used respectively the 1%, 2%, 5%, and 10% of the whole test set according to the chemical literature.
@@
As already mentioned, we compared the results of the multi-target classifier with another one developed for the same task by some of the authors. In that work molecules were described through a novel molecular embedding that we called EMBER. We carried out this comparison as in our previous work we showed that the EMBER classifier had SOTA performance.
In our results, both the accuracy and sensitivity values indicate that our multi-target network identifies correctly both the positive and negative samples. Particularly, sensitivity increased significantly with respect to the EMBER classifier where it ranged between 0.22 and 0.84. Now, sensitivity has an average value of 0.904, showing that it is much easier for the network to classify positive samples. The comparison between the two networks is summarised in Figure 3 showing the sensitivity for all targets. It is notable that there was a huge improvement in the performance related to sensitivity. Particularly, sensitivty on the CLK2 target raised from 0.22 to 0.938, as it is shown also in Table 1.
F1-Score values show that the network is able to overcome the dataset imbalance while simultaneously reducing the number of false negatives and false positives. The use of a properly balanced loss made by a weighted cross entropy and a soft F-score worked fine for the classification task. Finally, also the values for TP/P and EF are similar to the ones reported for the EMBER classifier.
@@ 3.2 Single-target classification
Classification performance in the CDK1-only network was assessed using the same metrics as described in Section 3.1 , and it is shown in Table 3 and 4.
@@
| CDK1 | 34/109 | 58/109 | 109/109 | 100/109 | 31.19 | 26.60 | 20 | 9.11 |
As can be observed, both Balanced Accuracy and Sensitivity are higher than those obtained on the same target in the multi-target classifier, while the performance in terms of both TP/P and EF remained consistent with the ones reported in Table 2.
@@ 3.3 Explainability results
As already detailed in Section 2.3, the main contribution of our work is the hierarchical Grad-CAM graph explainability procedure to devise relevant moieties at different scales in the molecule. Two visual examples of the results provided by our hierarchical graph explainer compared to the docking analysis carried out by an expert in Computational Chemistry are shown in Figure 4 and Figure 5. In such figures, two well known drugs acting as CDK1 inhibitors are reported: the Dinaciclib (PubChem ID: 46926350) and the Meridianin e (PubChem ID: 9995236). The color codes are the one already detailed in Section 2.3.
The explainability procedure was carried out on 19 CDK1 inhibitors, and the results have been compared to either the docking results for the same inhibitors or experimental crystal structures of co-crystallised ligands, when available. The complete visual results for all the drugs are reported in the Supplementary Materials. Our methodology is consistent with what was observed using experimental structures in 17 of 19 known drugs, and they are reported in Table 5 which lists the generic name of the drug, the PubChem ID, a binary assessment made by the expert on the goodness of prediction of the molecular features involved in the host-guest interaction between drug and target that we called Guest Feature Prediction (GFP) and the detailed accordance of the moieties coloring to the docking outcomes. Here, the expert verified if each piece of information gathered from the docking had been colored properly by our explainer that is with a label in accordance with the docking procedure. We reported such information in terms of single statements that have been red colored in case the explainer missed the correct label.
We performed also an ablation study to confirm our intuitions on the explainer procedure that is the need of a hierarchy of explainers acting at different neighborhood size, and the removal of the ReLU function in each Grad-CAM explainer to take into account also negative contributions to the inhibitory action of the ligand under investigation. The results are reported in Table 6, and they show clearly that the number of missed labels increases significantly both when we use a single layer explainer and if we leave the ReLU in the Grad-CAM output. Figure 6 reports the outputs provided by each layer in the hierarchy for the Dinaciclib. Interestingly, every layer of the explainer provides an important contribution to the final molecular interpretation. Indeed, the figure shows that the layer 3 explainer (the one considering the 3-neighborhood of each atom) is capable to catch important features as the H-bond donor N6, that are useful for the ligand protein complex stabilisation. Such a crucial information is then lost in the layer 7 explainer where molecular substructures are considered. On the other hand, the highest level of molecular interpretation is crucial to catch all the aromatic and lipophilic moieties that could be lost in the lower levels of the hierarchy.
| Drug | Pubchem ID | GFP | Accordance to docking | ||||
@@
| C26H31N7O* | 46916588 | yes |
*Molecular formula of 2-[(2S)-1-[8-ethyl-4-[(4-pyridin-2-ylphenyl)methylamino]pyrazolo[1,5-a][1,3,5]triazin-2-yl]piperidin-2-yl]ethanol
| Drug | Level 3 | Level 7 | Accordance to docking w/ ReLU | Our results (w/o ReLU) | ||||||||||||||||
@@
*Molecular formula of 2-[(2S)-1-[8-ethyl-4-[(4-pyridin-2-ylphenyl)methylamino]pyrazolo[1,5-a][1,3,5]triazin-2-yl]piperidin-2-yl]ethanol